# Subtype-Aware Registration of Longitudinal Electronic Health Records

### v1

@@ Abstract
Electronic Health Records (EHRs) contain extensive patient information that can inform downstream clinical decisions, such as mortality prediction, disease phenotyping, and disease onset prediction. A key challenge in EHR data analysis is the temporal gap between when a condition is first recorded and its actual onset time. Such timeline misalignment can lead to artificially distinct biomarker trends among patients with similar disease progression, undermining the reliability of downstream analysis and complicating tasks like disease subtyping. To address this challenge, we provide a subtype-aware timeline registration method that leverages data projection and discrete optimization to simultaneously correct timeline misalignment and improve disease subtyping. Through simulation and real-world data analyses, we demonstrate that the proposed method effectively aligns distorted observed records with the true disease progression patterns, enhancing subtyping clarity and improving performance in downstream clinical analyses.
@@ 1 Introduction
Electronic Health Records (EHR) are digitized collections of patient hospital information, encompassing data such as demographic, lab measurements, diagnoses, and medications. Data collected from intensive care unit (ICU) patients are of particular interest in clinical research due to their high frequency and timely collection. These data contain crucial information for executing timely disease interventions, managing patient care, and facilitating various downstream clinical decisions [1, 2, 3].
However, a significant challenge in utilizing EHR data lies in the temporal gap between when a condition is first recorded and its actual onset time. For example, a patient with pneumonia might be admitted to the hospital at the congestion stage (stage one) [4], while another patient might delay hospital visitation until a later stage due to factors like insurance coverage or personal health beliefs [5], resulting in unobserved data for the initial stages of disease progression. Conventional analyses on EHR data often overlook this misalignment of timelines, potentially leading to biased modeling results. Therefore, it is crucial to address this issue by develping methods that can effectively align observed EHR timelines with the true underlying disease progression.
@@ 1.1 Registration in the Literature
This need for timeline alignment is closely related to the problem of curve registration in functional data analysis, which seeks to align curves so that key features are properly synchronized across subjects [6, 7, 8]. In engineering, this problem is often referred to as time warping [9, 10]. Early methods aligned curves to a single baseline, using approaches like self-modeling nonlinear regression [11, 12] or template selection through functional principal component analysis (FPCA) [13], e.g., [14]. Subsequent work has explored a variety of alignment strategies. For instance, [15] proposed equating curve “moments” to capture local and global features, while [16] introduced a similarity index to iteratively register functional data. Other studies have applied Fisher-Rao metric-based approaches [17, 18, 19], computationally efficient techniques tailored for exponential family functional data [20, 21, 22], and pairwise warping strategies [23]. These approaches have been applied to a range of applications, including house price modeling [24], accelerometric activity data [25], and speech analysis [26]. However, since EHR data are often highly heterogeneous, sparse, noisy, and irregular, traditional curve registration methods are often not directly applicable.
@@
Furthermore, some studies have explored combining curve registration with clustering to account for the latent subgroups among subjects. [27] used an EM algorithm with B-splines, while [28] employed a Bayesian framework that accommodates flexible warping functions. Yet, these methods often rely on specific data distributions that may not align with the complexities of EHR datasets.
@@
To the best of our knowledge, [29] made a first attempt to address EHR-specific registration. They proposed aligning EHR trajectories by estimating time shifts to a common disease progression template. They assumed that all patients with a disease of interest share a common disease progression template and proposed a registration method using alternating optimization to align observed EHR data to this shared template. Despite its novelty, this approach has several limitations: (1) the reliance on linear interpolation and extrapolation for imputing missing data can lead to inaccurate estimations, particularly in sparsely observed EHR datasets; (2) the high computational burden of timeline registration for imputed curves becomes impractical for large patient cohorts; and (3) the assumption of a single progression template fails to account for the significant heterogeneity in diseases, where patients may follow distinct progression patterns.
@@
In summary, while various registration and clustering methods exist, they remain constrained by assumptions unsuitable for EHR data’s sporadic and heterogeneous nature. These challenges necessitate the development of a new timeline registration method that effectively address the distinct complexities of EHR datasets, including data sparsity, diverse disease subtypes, and large patient populations.
@@ 1.2 Our Contributions
To address the limitations mentioned above, this paper introduces an efficient and effective subtype-aware registration method tailored for electronic health records data. Our approach recognizes that patients can be grouped into subtypes, each potentially following a different disease progression pattern. Specifically, we consider an EHR dataset comprising $N$ subjects, where each subject $i$ has longitudinal data $\{(t_{ik},X_{i}(t_{ik}))\mid k=1,2,\ldots,T_{i}\}$. Here, $t_{ik}$ denotes the observation times within the domain of interest $[T_{\min},T_{\max}]$, $X_{i}(t_{ik})$ represents the corresponding observations, and $T_{i}$ is the number of observations for subject $i$. The observation times and the number of observations vary significantly across subjects. See the left panel of Figure 1 for an illustrative example on trajectories. The subjects belong to multiple disjoint subtypes $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$, where each subtype $\mathcal{P}_{k}$ contains patients with similar disease progression patterns.
Our goal is to align the timelines of subjects within each subtype more closely to each other by estimating individual time shifts $\tau_{i}$ and clustering labels $\pi_{i}$ such that the shifted data $X_{i}(t-\tau_{i})$ align optimally for all $i$ in subtype $\mathcal{P}_{k}$. Recognizing that outliers commonly occur in the electronic health records [30], we also aim to ensure robustness against them in our procedure.
To achieve this, we transform the observed longitudinal data and all possible time shifts using cubic B-spline regression, reducing the data’s dimensionality and enabling accelerated computation. We then cluster the data in the transformed space and iteratively update the clusters by selecting the time shifts that minimize the distance to cluster centroids. Each cluster in the transformed space is considered a potential subtype of the disease of interest. By focusing on minimizing intra-cluster variance, our method effectively and robustly captures the heterogeneity of disease progression patterns while maintaining robustness to outliers. In summary, our proposed method overcomes the limitations of existing approaches by incorporating subtype awareness, improving computational efficiency, and enhancing robustness to data sparsity and outliers. We present a simulation example in Figure 1 to demonstrate the effectiveness of timeline registration. The left panel shows the observed data and the right panel shows the unobserved underlying patterns, which include two unknown subtypes of trajectories marked in blue and red. The recovered trajectories after applying the proposed timeline registration are displayed in the center panel of Figure 1. Further discussions on the simulation studies are deferred to Section 3.
We evaluate our algorithm through extensive numerical studies. In six simulated datasets, our algorithm successfully realigned the observed data to the ground truth with less than one day of error in over 90% subjects. This realignment resulted in an enhanced separation between patient clusters, as measured by the Silhouette coefficient, from both data-driven and clinical perspectives. Furthermore, the realigned EHR data improved the predictive accuracy for severe AKI patients by 3–5%, demonstrating the effectiveness of the registered data in supporting accurate prediction modeling.
(a) Step 1: Time Series Transformation Using Cubic B-spline Regression
(b) Step 2: Iterations
(c) Step 3: Finalization
Figure 2: This figure illustrates subtype-aware registration using longitudinal EHR data. Black lines and dots represent the current curves, observations, and their low-dimensional embeddings. Red lines and dots indicate curves and observations that include potential time shifts under consideration. Blue dots represent potential outliers that were excluded from centroid evaluation for the time being.
@@ 2 Method for Subtype-Aware Registration of EHR
[…]
Next, we propose the EHR registration algorithm shown in Figure 2, which consists of three main steps, detailed in the following subsections. The pseudo-code for each step is provided in Algorithms 1, 2, and 3.
@@
In this step, we apply cubic B-spline regression [31, 32] to perform time series transformation. This technique models longitudinal data with flexible and smooth approximations, yielding a low-dimensional representation.
Specifically, we introduce $q$ internal knots, denoted as $\{\tau_{1},\ldots,\tau_{q}\}$, within the overall time domain $[T_{\min},T_{\max}]$. Based on these knots, we construct $p=q+3$ basis functions for cubic splines $B_{j,3}(t)$ recursively as follows.
@@
$B_{j,0}(t)=\left\{\begin{array}[]{ll}1&\text{if }t\in[\tau_{j},\tau_{j+1}),\\ 0&\text{otherwise}.\end{array}\right.$
Then, for $m=1,2,3$, the higher-degree basis functions are constructed recursively:
$B_{j,m}(t)=\frac{t-\tau_{j}}{\tau_{j+m}-\tau_{j}}B_{j,m-1}(t)+\frac{\tau_{j+m+1}-t}{\tau_{j+m+1}-\tau_{j+1}}B_{j+1,m-1}(t).$
For more details about B-splines, readers are referred to [33].
We perform B-spline regression using ridge regression (which provides more robustness) to fit both the original and time-shifted observations for each subject $i=1,\ldots,N$. Specifically, we pre-specify a set of potential shift positions $\Lambda=\{l_{0},l_{1},\ldots,l_{m}\}$, where $l_{0}=0$ means no shift. For each $i=1,\ldots,N$ and $l\in\Lambda$, we perform ridge regression $X_{i}(t-l)\sim B_{j,3}(t)$ by evaluating the following estimator:
$\widehat{\boldsymbol{\beta}}_{i,l}=\mathop{\rm arg\min}_{\boldsymbol{\beta}_{i,l}}\left\{\sum_{\begin{subarray}{c}k=1\\ T_{\min}\leq t_{ik}-l\leq T_{\max}\end{subarray}}^{T_{i}}\left(X_{i}(t_{ik}-l)-\sum_{j=1}^{p}\beta_{ij}B_{j,3}(t_{ik})\right)^{2}+\lambda\sum_{j=1}^{p}\beta_{ij}^{2}\right\},$
where $\lambda$ is the tuning parameter that controls the penalization level, and $\widehat{\boldsymbol{\beta}}_{i,l}=(\widehat{\beta}_{i1},\widehat{\beta}_{i2},\ldots,\widehat{\beta}_{ip})$ is the vector of estimated coefficients. In this way, the underlying structure and dynamics of the longitudinal data are reduced to $p$ coefficients. This cubic B-spline regression yields the following low-dimensional embedding:
$\{(t_{ik},X_{i}(t_{ik}))\mid k=1,2,\ldots,T_{i}\}\quad\longrightarrow\quad\widehat{\boldsymbol{\beta}}_{i,l}.$
@@
We set the initial time shifts as $\delta_{i}^{(0)}=0$ for all $i=1,\ldots,N$. Next, we run the following steps for $t=1,2,\ldots$:
@@
    Clustering with Optimal Number of Clusters. For each candidate number of clusters $k\in\{2,\ldots,M\}$, we conduct clustering (e.g., $k$-means, $k$-medoids) with $k$ clusters on the low-dimensional representations with time shifts $\delta_{i}^{(t-1)}$: $\{\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}\}_{i=1}^{N}.$ When $t=1$, the data is the original longitudinal dataset without time shifts.
    We determine the number of clusters $K^{(t)}$ by selecting the $k$ that maximizes the average Silhouette coefficient [34]:
    $K^{(t)}=\mathop{\rm arg\max}_{k\in\{2,\ldots,M\}}s(k),\quad s(k)=\frac{1}{N}\sum_{i=1}^{N}s(i,k),$ (1)
    where $s(i,k)$ is the Silhouette coefficient for subject $i$ when the number of clusters is $k$.
    This results in $K^{(t)}$ clusters of the subjects: $\{\mathcal{C}_{1}^{(t)},\ldots,\mathcal{C}_{K^{(t)}}^{(t)}\}\subseteq[N]$, where $\mathcal{C}_{i}^{(t)}\cap\mathcal{C}_{j}^{(t)}=\emptyset$ for all $i\neq j$.
@@
    Iterative Updates. For each cluster $k=1,\ldots,K^{(t)}$, we perform:
    • Compute the centroid of cluster $\mathcal{C}_{k}^{(t)}$:
      $\bar{\boldsymbol{\beta}}_{k}^{(t)}=\frac{1}{|\mathcal{C}_{k}^{(t)}|}\sum_{i\in\mathcal{C}_{k}^{(t)}}\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}.$
@@
      $d_{i}=\left\|\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}-\bar{\boldsymbol{\beta}}_{k}^{(t)}\right\|,\quad\forall i\in\mathcal{C}_{k}^{(t)}.$
    • Select the subset $\mathcal{S}_{k}^{(t)}$ of points corresponding to the smallest $\left\lceil\alpha|\mathcal{C}_{k}^{(t)}|\right\rceil$ distances:
      $\mathcal{S}_{k}^{(t)}=\left\{i\in\mathcal{C}_{k}^{(t)}\mid d_{i}\text{ is among the smallest }\left\lceil\alpha|\mathcal{C}_{k}^{(t)}|\right\rceil\text{values}\right\},$
@@
      $\bar{\boldsymbol{\gamma}}_{k}^{(t)}=\frac{1}{|\mathcal{S}_{k}^{(t)}|}\sum_{i\in\mathcal{S}_{k}^{(t)}}\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}.$
      This step removes potential outliers ($\mathcal{C}_{k}^{(t)}\setminus\mathcal{S}_{k}^{(t)}$) from the centroid computation, enhancing robustness.
    • Update the time shifts for each subject $i\in\mathcal{C}_{k}^{(t)}$ by finding the shift $\delta_{i}^{(t)}\in\Lambda$ that minimizes the distance to the updated centroid:
      $\delta_{i}^{(t)}=\mathop{\rm arg\min}_{\delta\in\Lambda}\left\|\mathbf{\Omega}_{i,\delta}-\bar{\boldsymbol{\gamma}}_{k}^{(t)}\right\|^{2}.$

### v2

@@ Abstract
Electronic Health Records (EHRs) contain extensive patient information that can inform downstream clinical decisions, such as mortality prediction, disease phenotyping, and disease onset prediction. A key challenge in EHR data analysis is the temporal gap between when a condition is first recorded and its actual onset time. Such timeline misalignment can lead to artificially distinct biomarker trends among patients with similar disease progression, undermining the reliability of downstream analyses and complicating tasks such as disease subtyping and outcome prediction. To address this challenge, we provide a subtype-aware timeline registration method that leverages data projection and discrete optimization to correct timeline misalignment. Through simulation and real-world data analyses, we demonstrate that the proposed method effectively aligns distorted observed records with the true disease progression patterns, enhancing subtyping clarity and improving performance in downstream clinical analyses.
Abstract
We provide simulation studies in these supplementary materials.
@@ 1 Introduction
Electronic Health Records (EHR) are digitized collections of patient hospital information, encompassing data such as demographics, lab measurements, diagnoses, and medications. Data collected from intensive care unit (ICU) patients are of particular interest in clinical research due to their high frequency and timely collection. These data contain crucial information for executing timely disease interventions, managing patient care, and facilitating various downstream clinical decisions Shickel2017DeepEA ; Yadav2017MiningEH ; Shillan2019UseOM .
However, a significant challenge in utilizing EHR data lies in the temporal gap between when a condition is first recorded and its actual onset time. For example, a patient with pneumonia might be admitted to the hospital at the congestion stage (stage one) Torres2021Pneumonia , while another patient might delay hospital visitation until a later stage due to factors like insurance coverage or personal health beliefs Irlmeier2022Cox , resulting in unobserved data for the initial stages of disease progression. Conventional analyses on EHR data often overlook this misalignment of timelines, potentially leading to biased modeling results. Therefore, it is crucial to address this issue by developing methods that can effectively align observed EHR timelines with the true underlying disease progression.
@@ 1.1 Registration in the Literature
This need for timeline alignment is closely related to the problem of curve registration in functional data analysis, which seeks to align curves so that key features are properly synchronized across subjects wang2016functional ; marron2015functional ; Ramsay2002CurveR . In engineering, this problem is often referred to as time warping sakoe1978dynamic ; wang1999synchronizing . Early methods aligned curves to a single baseline, using approaches like self-modeling nonlinear regression lawton1971self ; kneip1988convergence or template selection through functional principal component analysis (FPCA) yao2005functional , e.g., Kneip2008CombiningRA . Subsequent work has explored a variety of alignment strategies. For instance, James james2007curve proposed equating curve “moments” to capture local and global features, while Sangalli et al. sangalli2009case introduced a similarity index to iteratively register functional data. Other studies have applied Fisher-Rao metric-based approaches srivastava2011registration ; tucker2013generative ; wu2014analysis , computationally efficient techniques tailored for exponential family functional data wrobel2019registration ; wrobel2018register ; wrobel2021registr , and pairwise warping strategies tang2008pairwise . These approaches have been applied to a range of applications, including house price modeling peng2014time , accelerometric activity data mcdonnell2022registration , and speech analysis hadjipantelis2015unifying . However, since EHR data are often highly heterogeneous, sparse, noisy, and irregular, traditional curve registration methods are often not directly applicable.
@@
Furthermore, some studies have explored combining curve registration with clustering to account for the latent subgroups among subjects. Liu et al. liu2009simultaneous used an EM algorithm with B-splines, while Wu et al. wu2016bayesian employed a Bayesian framework that accommodates flexible warping functions. Yet, these methods often rely on specific data distributions that may not align with the complexities of EHR datasets.
@@
To the best of our knowledge, Jiang et al. jiang2023timeline made a first attempt to address EHR-specific registration. They proposed aligning EHR trajectories by estimating time shifts to a common disease progression template. They assumed that all patients with a disease of interest share a common disease progression template and proposed a registration method using alternating optimization to align observed EHR data to this shared template. Despite its novelty, this approach has several limitations: (i) the reliance on linear interpolation and extrapolation for imputing missing data can lead to inaccurate estimations, particularly in sparsely observed EHR datasets; (ii) the high computational burden of timeline registration for imputed curves becomes impractical for large patient cohorts; and (iii) the assumption of a single progression template fails to account for the significant heterogeneity in diseases, where patients may follow distinct progression patterns.
@@
In summary, while various registration and clustering methods exist, they remain constrained by assumptions unsuitable for EHR data’s sporadic and heterogeneous nature. These challenges necessitate the development of a new timeline registration method that effectively addresses the distinct complexities of EHR datasets, including data sparsity, diverse disease subtypes, and large patient populations.
@@ 1.2 Our Contributions
In this paper, we develop a subtype-aware timeline registration framework for longitudinal EHR data, with a focus on ICU AKI trajectories. Our registration framework simultaneously aligns patient timelines and discovers latent progression subtypes. By allowing different clusters of patients to follow distinct temporal patterns, our method moves beyond single-template registration and better reflects the clinical heterogeneity of ICU populations. In the proposed registration framework, we design a three-step algorithm that first embeds irregular trajectories into a low-dimensional B-spline representation, then iteratively alternates between clustering and optimizing subject-specific time shifts, and finally refines assignments while trimming potential outliers. This design yields substantial computational savings and robustness to noisy measurements and aberrant trajectories, making it suitable for large-scale EHR databases.
Through extensive simulations, we show that the proposed method can recover latent time shifts with small error across a wide range of trajectory shapes, mean structures, and progression speeds, with runtimes under a few minutes. In real data from the MIMIC-IV ICU AKI cohort, our registration improves phenotype separation, yields clinically interpretable AKI subtypes, and increases severe AKI prediction performance by 3–5% compared to unregistered data and an existing EHR registration baseline. The corresponding R package implementing the proposed registration method is publicly available at https://github.com/LeoGai/subtypeReg, and the source code used to reproduce the analyses in this paper can be found at https://github.com/LeoGai/EHR-registration-subtype.
1.3 Organization of the paper
The rest of the paper is organized as follows. In Section 2, we describe the ICU EHR database, the clinical question, and our formal problem setup. Section 3 presents the proposed subtype-aware registration algorithm and its computational properties. Section 4 presents a comprehensive analysis of AKI trajectories in the MIMIC-IV ICU cohort, including unsupervised phenotyping, clinical evaluation, and supervised prediction. We conclude with discussions of limitations and future directions in Section 5. Supplementary Section A reports extensive simulation results, and Supplementary Section C provides an analysis of the computational and space complexities.
2 EHR Database and Problem Formulation
2.1 ICU EHR Database and Clinical question
We analyze ICU EHR data from the MIMIC-IV database, which contains detailed clinical information on over 40,000 ICU stays at the Beth Israel Deaconess Medical Center between 2008 and 2019 Goldberger2000PhysionetCO ; Johnson2023MIMICIVAF ; johnson2023mimicivdemo . For this study, we focus on patients with acute kidney injury (AKI), a common clinical condition characterized by an abrupt decline in kidney function, resulting from various causes such as decreased renal perfusion, direct nephron injury, or urinary tract obstruction Makris2016AcuteKI . AKI can develop rapidly within hours or days and is associated with significant morbidity and high mortality rates Rewa2014AcuteKI . This syndrome is highly heterogeneous, and the complexity of its pathophysiology poses significant challenges in developing effective patient treatment plans Rewa2014AcuteKI .
We identify ICU stays with ICD codes indicating acute kidney failure and text descriptions containing the phrase “acute kidney injury.” We exclude patients with burns, end-stage renal disease, or renal dialysis during their ICU stay, as well as those with an estimated glomerular filtration rate (eGFR) below 15 mL/min/1.73 m2 to avoid including end-stage renal disease song2020cross . For data quality, we remove ICU stays with fewer than 15 recorded serum creatinine (SCr) measurements in the first 21 days after ICU admission, and those with two or fewer SCr measurements in each of the intervals 0–7, 7–14, and 14–21 days.
Serum creatinine is the primary laboratory marker used to diagnose and stage AKI and to guide treatment decisions Khwaja2012KDIGOCP . However, the first recorded SCr in the ICU does not generally correspond to the same disease stage across patients. Some patients are admitted early in their AKI course, whereas others arrive after several days of unobserved deterioration. Our central clinical questions are:
  1. 1.
    (Latent timeline recovery) Can we recover a latent disease progression timeline from irregular and heterogeneous ICU EHR trajectories?
  2. 2.
    (Cross-patient registration and subtype discovery) Can these latent timelines be registered across patients to identify clinically meaningful AKI subtypes?
  3. 3.
    (Downstream utility) Can the recovered and registered disease timelines improve downstream clinical analyses?
Answering these questions requires an EHR-tailored registration method that respects the sparsity, irregularity, and heterogeneity of ICU data, and scales to large cohorts.
2.2 Timeline misalignment and statistical formulation
We consider an EHR dataset comprising $N$ ICU stays (subjects), indexed by $i=1,\ldots,N$. For each subject $i$, let $\{(t_{ir},X_{i}(t_{ir})):r=1,\ldots,T_{i}\}$ denote their longitudinal SCr measurements, where $t_{ir}$ is the time (in days) since ICU admission and $T_{i}$ is the number of observations. The observation times $t_{ir}$ lie in a common window $T_{\min},T_{\max}]$ but are irregular and subject-specific; both the number and timing of measurements vary substantially across patients.
We postulate that patients can be grouped into latent AKI subtypes, with each subtype following a distinct underlying disease progression pattern. Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$ denote $K$ disjoint subtypes, and let $\pi_{i}\in\{1,\ldots,K\}$ be the subtype label for subject $i$. For a given subtype $k$, we assume there exists an underlying “intrinsic” disease timeline such that trajectories within $\mathcal{P}_{k}$ are approximately aligned after subject-specific time shifts $\delta_{i}$. Intuitively, $\delta_{i}$ represents the delay between true disease onset and ICU admission for subject $i$.
Our goal is to estimate both the subtype labels $\{\pi_{i}\}$ and the time shifts $\{\delta_{i}\}$ from the observed, irregularly sampled trajectories $\{X_{i}(t_{ir})\}$, such that the shifted trajectories $\{(t_{ir}+\delta_{i},\;X_{i}(t_{ir}))\}_{r=1}^{T_{i}}$ are well aligned within each subtype. At the same time, we must be robust to outliers and atypical trajectories, which are common in ICU EHR data hauskrecht2013outlier . Formally, we aim to find clustering and shifts that minimize within-subtype dispersion of the aligned trajectories while downweighting or trimming potential outliers.
@@ 3 Method for Subtype-Aware Registration of EHR
To address this problem, we develop a scalable, subtype-aware registration algorithm tailored to longitudinal EHR data. The algorithm proceeds in three steps (illustrated in Figure 1):
  1. 1.
    Spline-based time series embedding. We first map each irregular trajectory, under a grid of candidate time shifts, to a low-dimensional feature representation using cubic B-spline regression with ridge regularization. This yields a compact representation of each subject–shift combination, greatly reducing the dimensionality and enabling efficient comparison of trajectories with different sampling patterns.
  2. 2.
    Iterative clustering and time-shift refinement. Starting from zero shifts, we repeatedly cluster the current embeddings (using $k$-medoids or $k$-means) and update subject-specific time shifts to bring each trajectory closer to its cluster centroid in the embedding space. At each iteration, we also compute trimmed centroids based on the most central trajectories within each cluster, which improves robustness to outliers and poorly aligned subjects.
  3. 3.
    Final assignment of outlying trajectories. After convergence of the iterative procedure, we reassign previously trimmed or outlying trajectories by jointly optimizing over candidate time shifts and cluster centroids. The final output is a set of estimated shifts $\{\widehat{\delta}_{i}\}$ and subtype labels $\{\widehat{\pi}_{i}\}$ that define aligned AKI progression trajectories for downstream analysis.
Figure 1: Illustration of subtype-aware registration using longitudinal EHR data.
Next, we describe the proposed registration algorithm in detail. The pseudo-code for each step is provided in Algorithms 13. Through this paper, we denote $N]=\{1,\ldots,N\}$ for any integer $N\geq 1$. Let $|\mathcal{C}|$ be the cardinality of any set $\mathcal{C}$. The Euclidean norm or the $\ell_{2}$ norm of a vector $x$ is denoted as $\left\|\cdot\right\|$, i.e., $\|x\|=\sqrt{\sum_{i}x_{i}^{2}}$. Then, $\left\|\mathbf{u}-\mathbf{v}\right\|$ represents the Euclidean distance between two vectors $\mathbf{u}$ and $\mathbf{v}$.
@@
In this step, we apply cubic B-spline regression (wahba1990spline, ; hastie2009elements, ) to perform time series transformation. This technique models longitudinal data with flexible and smooth approximations, yielding a low-dimensional representation.
Specifically, we introduce $q+8$ knots within the overall time domain $T_{\min},T_{\max}]$, denoted as $T_{\min}\leq\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}\leq\cdots\leq\kappa_{q+5}=\kappa_{q+6}=\kappa_{q+7}=\kappa_{q+8}\leq T_{\max}$. Based on these knots, we construct $p=q+4$ basis functions for cubic splines $B_{j,3}(t)$ recursively as follows.
@@
$B_{j,0}(t)=\left\{\begin{array}[]{ll}1&\text{if }t\in[\kappa_{j},\kappa_{j+1}),\\ 0&\text{otherwise}.\end{array}\right$.
Then, for $d=1,2,3$, the higher-degree basis functions are constructed recursively:
$B_{j,d}(t)=\frac{t-\kappa_{j}}{\kappa_{j+d}-\kappa_{j}}B_{j,d-1}(t)+\frac{\kappa_{j+d+1}-t}{\kappa_{j+d+1}-\kappa_{j+1}}B_{j+1,d-1}(t).$
For more details about B-splines, readers are referred to wahba2014spline . Because the B-spline basis functions sum to one, including an intercept together with all $B_{j,3}(t)$ leads to linear dependence. We therefore include an intercept and drop one basis function $B_{1,3}$ to ensure identifiability. If the boundary knots $\kappa_{1},\kappa_{q+8}$ were fixed globally, shifting a patient’s timeline could leave sparse or no observations near the boundaries, so the fitted spline would be weakly constrained by the data there and could become unstable. To mitigate this instability in practice, we can also set the boundary knots to the minimum and maximum observed time points for each shifted trajectory, after applying the candidate time shifts.
We perform B-spline regression using ridge regression (which improves numerical stability) to fit both the original and time-shifted observations for each subject $i=1,\ldots,N$. Specifically, we pre-specify a set of potential shift positions $\Lambda=\{l_{0},l_{1},\ldots,l_{m}\}$, where $l_{0}=0$ means no shift. For each $i=1,\ldots,N$ and $l\in\Lambda$, we perform ridge regression $X_{i}(t)\sim B_{j,3}(t+l)$ by evaluating the following estimator:
$\widehat{\boldsymbol{\beta}}_{i,l}=\mathop{\rm arg\min}_{\boldsymbol{\beta}_{i,l}}\left\{\sum_{\begin{subarray}{c}r=1\\ T_{\min}\leq t_{ir}+l\leq T_{\max}\end{subarray}}^{T_{i}}\left(X_{i}(t_{ir})-\beta_{i,l,0}-\sum_{j=2}^{p}\beta_{i,l,j}B_{j,3}(t_{ir}+l)\right)^{2}+\lambda\sum_{j=2}^{p}\beta_{i,l,j}^{2}\right\},$
where $\lambda$ is the tuning parameter that controls the penalization level, and $\widehat{\boldsymbol{\beta}}_{i,l}=(\widehat{\beta}_{i,l,0},\widehat{\beta}_{i,l,2},\ldots,\widehat{\beta}_{i,l,p})$ is the vector of estimated coefficients. In this way, the underlying structure and dynamics of the longitudinal data are reduced to $p$ regression coefficients (including an intercept). This cubic B-spline regression yields the following low-dimensional embedding:
$\{(t_{ir},X_{i}(t_{ir}))\mid r=1,2,\ldots,T_{i}\}\quad\longrightarrow\quad\widehat{\boldsymbol{\beta}}_{i,l}.$
@@
We set the initial time shifts as $\delta_{i}^{(0)}=0$ for all $i=1,\ldots,N$. Next, we run the following steps for $h=1,2,\ldots$:
@@
    Clustering with Optimal Number of Clusters. For each candidate number of clusters $k\in\{2,\ldots,M\}$, we conduct clustering (e.g., $k$-means, $k$-medoids) with $k$ clusters on the low-dimensional representations with time shifts $\delta_{i}^{(h-1)}$: $\{\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}\}_{i=1}^{N}.$ When $h=1$, the data is the original longitudinal dataset without time shifts.
    We determine the number of clusters $K^{(h)}$ by selecting the $k$ that maximizes the average silhouette coefficient (rousseeuw1987silhouettes, ):
    $K^{(h)}=\mathop{\rm arg\max}_{k\in\{2,\ldots,M\}}s(k),\quad s(k)=\frac{1}{N}\sum_{i=1}^{N}s(i,k),$
    where $s(i,k)$ is the silhouette coefficient for subject $i$ when the number of clusters is $k$, defined as:
    $s(i,k)=\frac{b(i)-a(i)}{\max\{a(i),b(i)\}}.$
    Here, $a(i)$ is the average distance of subject $i$ to all other subjects within the same cluster and $b(i)$ is the lowest average distance of subject $i$ to any other cluster to which $i$ does not belong, calculated as:
    $a(i)=\frac{1}{|\mathcal{C}_{w}|-1}\sum_{j\in\mathcal{C}_{w},j\neq i}\|\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}-\mathbf{\Omega}_{j,\delta_{j}^{(h-1)}}\|,\quad b(i)=\min_{w^{\prime}\neq w}\frac{1}{|\mathcal{C}_{w^{\prime}}|}\sum_{j\in\mathcal{C}_{w^{\prime}}}\|\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}-\mathbf{\Omega}_{j,\delta_{j}^{(h-1)}}\|,$
    where $\mathcal{C}_{w}$ is the cluster to which subject $i$ belongs to.
    This results in $K^{(h)}$ clusters of the subjects: $\{\mathcal{C}_{1}^{(h)},\ldots,\mathcal{C}_{K^{(h)}}^{(h)}\}\subseteq[N]$, where $\mathcal{C}_{i}^{(h)}\cap\mathcal{C}_{j}^{(h)}=\emptyset$ for all $i\neq j$.
@@
    Iterative Updates. For each cluster $k=1,\ldots,K^{(h)}$, we perform:
    • Compute the centroid of cluster $\mathcal{C}_{k}^{(h)}$:
      $\bar{\boldsymbol{\beta}}_{k}^{(h)}=\frac{1}{|\mathcal{C}_{k}^{(h)}|}\sum_{i\in\mathcal{C}_{k}^{(h)}}\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}.$
@@
      $d_{i}=\left\|\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}-\bar{\boldsymbol{\beta}}_{k}^{(h)}\right\|,\quad\forall i\in\mathcal{C}_{k}^{(h)}.$
    • Select the subset $\mathcal{S}_{k}^{(h)}$ of points corresponding to the smallest $\left\lceil\alpha|\mathcal{C}_{k}^{(h)}|\right\rceil$ distances:
      $\mathcal{S}_{k}^{(h)}=\left\{i\in\mathcal{C}_{k}^{(h)}\mid d_{i}\text{ is among the smallest }\left\lceil\alpha|\mathcal{C}_{k}^{(h)}|\right\rceil\text{ values}\right\},$
@@
      $\bar{\boldsymbol{\gamma}}_{k}^{(h)}=\frac{1}{|\mathcal{S}_{k}^{(h)}|}\sum_{i\in\mathcal{S}_{k}^{(h)}}\mathbf{\Omega}_{i,\delta_{i}^{(h-1)}}.$
      This step removes potential outliers ($\mathcal{C}_{k}^{(h)}\setminus\mathcal{S}_{k}^{(h)}$) from the centroid computation, enhancing robustness.
    • Update the time shifts for each subject $i\in\mathcal{C}_{k}^{(h)}$ by finding the shift $\delta_{i}^{(h)}\in\Lambda$ that minimizes the distance to the updated centroid:
      $\delta_{i}^{(h)}=\mathop{\rm arg\min}_{\delta\in\Lambda}\left\|\mathbf{\Omega}_{i,\delta}-\bar{\boldsymbol{\gamma}}_{k}^{(h)}\right\|^{2}.$